OpenStudy (anonymous):

solve for x in the given interval sinxtan3x=0(-pi,0)

OpenStudy (anonymous):

\[sinxtan3x=0(-\pi,0)?\]

OpenStudy (anonymous):

\[well, x = 2 \pi n ~~ 6.2832 n and n element Z\]

OpenStudy (anonymous):

\[x = 0\]

OpenStudy (anonymous):

how did you get the 0? I got the -pi

OpenStudy (anonymous):

also there is a couple more answers

OpenStudy (anonymous):

sin x tan x=0 (-pi,0) either sin x=0=sin npi ,where n is an integer. x=-pi,x=0 if (-pi,0) is an open interval, then there is no solution. or tan x=0=tan npi,where n is an integer. x=-pi,0 in the closed interval ( pi,0) in the open interval it has no solution.

OpenStudy (anonymous):

correction tan 3x=0=tan npi ,where n is an integer. 3x=n pi,x=npi/3 (put n/3=-1,n=-3) for n=-3,-2,-1,0 x=-pi,-2pi/3,-pi/3,0

OpenStudy (anonymous):

for open interval x=-2pi/3,-pi/3 are the solutions.

OpenStudy (anonymous):

how are you getting the npi thing? what is that? and what values do you plug into that